OpenStudy (anonymous):

The time required for an automotive center to complete an oil change service on an automobile approximately follows a normal distribution, with the mean of 17 minutes and a standard deviation of 2 minutes. a) The automotive center guarantees customers that the service will take no longer than 20 minutes. If it does take longer, the customer will receive the service for half-price. What percent of customers receive the service for half-price? b) If the automotive center does not want to give the discount to more than 2% of its customers, how long should it make the guaranteed time limit?

OpenStudy (anonymous):

Hi, Those customers who receive half-price are those where the service takes more than 20 minutes. Since the mean is 17 minutes, the customers who receives half-price are those where the service takes 20 - 17 = 3 minutes above the mean.

OpenStudy (anonymous):

When dealing with these kind of problems, the thing is all about converting units. You need to convert those 3 units ''of minutes'' in units of standard deviation.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

You are given the following information:

OpenStudy (anonymous):

\(\large{\mu = 17\ minutes}\) \(\large{\sigma = 2 \frac{minutes}{standard\ deviation}}\) <- It can be read as 'minutes per standard deviation'. \(\large{X = 20\ minutes}\)

OpenStudy (anonymous):

A unit conversion (for instance, from inches to meters) is all about cancelling units.

OpenStudy (anonymous):

We want to find how many standard deviations away those 3 minutes lie above the mean. Cancelling unit of minutes and leaving units of standard deviations is needed. Do you guess how it can be done?

OpenStudy (anonymous):

I'm not sure. Cross multiplication?

OpenStudy (anonymous):

Yep. Let me help you: \(\large{\frac{3\ minutes}{\left(2 \frac{minutes}{standard\ deviation}\right)} = \frac{3}{2} minutes * \frac{standard\ deviation}{minutes}}\)

OpenStudy (anonymous):

minutes cancels out and we are left with 3/2 standard deviations (sd) = 1.5 (sd)

OpenStudy (anonymous):

We are almost done with a), because you are asked to find the percentage of customer where the service takes more than 3 minutes. We know that more than 3 minutes means more than 1.5 sd above the mean.

OpenStudy (anonymous):

To find this out, you need a z-table or z-calculator. Do you know how to do it?

OpenStudy (anonymous):

I have a z-table. I was curious, how did you get 1.5 sd?

OpenStudy (anonymous):

Take a look the last expression I wrote in Latex code. 3/2 = 1.5 and minutes cancel each other out, leaving just the units of standard deviation. The 1.5 sd (:

OpenStudy (anonymous):

I understand. What would I have to look for on my z table?

OpenStudy (anonymous):

Since you want to find the probability of getting a value (an oil change service) greater than 1.5 sd from the mean, you need to look at the proportion that lies 1.5 sd toward the tail.

OpenStudy (anonymous):

I'm not sure how to read this :[

OpenStudy (anonymous):

Usually a z-table has 4 columns. What we've found, 1.5 sd is called the z-score.

OpenStudy (anonymous):

Take a look at the image below.

OpenStudy (anonymous):

Proportion in tail is what we are looking for. It just means the proportion above 1.5sd and it goes all the way to the upper tail.

OpenStudy (anonymous):

Does it seem too difficult?

OpenStudy (anonymous):

Yes. I'm confused.

OpenStudy (anonymous):

Notice the process we followed: First we find the difference of the value given (X) from the mean. Then we convert from those units to unit of standar deviation. This is generally written as: \(\large{\frac{X-\mu}{\sigma}}\) Then we get the z-score. And finally we find the probability. Having X and mean -> Difference -> unit conversion -> z-score -> probability.

OpenStudy (anonymous):

To solve b) you need to do the whole process in reverse (following the arrow from the right): Having mean but not X <- Difference <- unit conversion <- z-score <- probability